# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

### v2

@@ 4 Experiments
In the following sections, we examine whether combining LRRL with Adam or RMSProp —two widely used optimization algorithms— can enhance cumulative returns in deep RL algorithms. We compare LRRL against learning methods both with and without schedulers, using baseline algorithms provided in the Dopamine [9] framework. The experiments involve testing LRRL across various configurations and a selection of Atari games, reporting the mean and one-half standard deviation of returns over 5 independent runs. As testing the full Atari benchmark is extremely costly, we choose a representative subset of games guided by their baseline learning curves, which offer insights into reward sparsity, stochasticity, and exploration demands. Further details on the evaluation metrics and hyperparameters used in these experiments are provided in Appendices C and D. In addition, although our primary target is RL tasks, we also test LRRL using SGD to minimize stationary objective functions to provide a simpler evaluation, as performance in RL is highly multi-factorial.
4.1 Comparing LRRL with Standard Learning
In our first experiment, we consider a set of $5$ learning rates. We set different configurations of LRRL using these values and compare them against the DQN agent reaching the best performance, in terms of maximum average return, among the possible single rates. More specifically, the learning rates considered are
@@ 4.1 Comparing LRRL with Standard Learning
The whole set $\mathcal{K}(5)$ of rates is used by one LRRL version, while others are based on the $3$ lowest ($\mathcal{K}_{\text{lowest}}(3))$, $3$ middle ($\mathcal{K}_{\text{middle}}(3))$, $3$ highest ($\mathcal{K}_{\text{highest}}(3)$) values, and the last one ($\mathcal{K}_{\text{sparse}}(3)$) taking the lowest/middle/highest values. All experiments use the Adam optimizer, and we report the return based on the same number of environment iterations. The results in Figure 1 demonstrate that LRRL using the full set $\mathcal{K}(5)$ substantially outperforms the best-performing single learning rate in 2 out of 6 games (Seaquest and Breakout) while remaining competitive in the others. As shown in Figure 7 in Appendix B.1, the baseline DQN performance is highly sensitive to the learning rate choice, requiring extensive hyperparameter tuning to identify the optimal rate. In contrast, LRRL dynamically selects learning rates during training, avoiding the need for such tuning. Additionally, the $3$-arm bandit variants highlight that while LRRL can explore multiple learning rates within a single run, the specific set of rates available still impacts performance. Overall, these results illustrate the advantage of LRRL, which achieves similar or better performance than the best-tuned DQN baseline while eliminating the computational cost of manual learning rate tuning.
Figure 1: Comparison between LRRL and standard DQN learning: 5 variants of LRRL with different learning rate sets (provided in Appendix D) are tested again the DQN algorithm reaching best performance among the possible learning rates. The mean and one-half standard deviations over 5 runs are respectively represented by the curves and shaded areas.
Next, we illustrate how LRRL adapts during training in response to non-stationary bandit feedback as policy performance improves. Figure 2 shows the systematic sampling of pulled arms (learning rates) and corresponding returns over training steps from a single run of LRRL $\mathcal{K}(5)$. In most of the tested environments, LRRL behaves similarly to time-decay schedulers by selecting higher learning rates during the early stages of training, gradually shifting toward arms with lower rates as training progresses.
@@
4.2 Combining and Comparing Schedulers with LRRL
@@
In the following experiments, LRRL is combined with learning rate schedulers. Specifically, we employ schedulers with exponential decay rate of the form $\eta(n)=\eta_{0}\times e^{-d\,n}$, where $\eta_{0}$ is a fixed initial value (common to each scheduler and equal to $6.25\times 10^{-5}$ in our experiment), $d$ is the exponential decay rate and $n$ is the number of policy updates (i.e., of MAB rounds). We define a set of $3$ schedulers $\mathcal{K}_{s}$, where each arm represents a scheduler using a different decay rate $d=\{1,2,3\}\times 10^{-7}$, and compare the results of LRRL with each scheduler individually, using the Adam optimizer.
Figure 3 shows that LRRL combined with schedulers could once again substantially increase the final performance of 2 out of 6 tasks compared to using a fixed scheduler. In the remaining tasks, LRRL remains at least competitive, while for most of them, it shows a better jumpstart performance. The dashed black line represents the maximal average return achieved by Adam without learning rate decay, resulting in slightly worse performance compared to using schedulers, aligning with findings in previous work by Andrychowicz et al. [4], who linearly decay the learning rate to 00.
In all the experiments conducted so far, we have set the update window $\kappa$ parameter to 1, which allows the observation of a full episode before updating the learning rate. In Appendix B.3, we address how this value could influence the performance of LRRL by allowing the observation of more interactions.
@@
Figure 3: LRRL with arms that select learning rate schedulers: LRRL vs 3 individual schedulers ($d=1,2,3$). The dashed black line represents the max average return achieved by Adam with a constant learning rate set to $6.25\times 10^{-5}$.
4.3 Comparing RMSProp and Adam Optimizers
Another widely used optimizer for deep RL models is RMSProp, which, like Adam, features an adaptive learning rate mechanism. Adam builds upon RMSProp by retaining exponential moving averages to give more weight to recent gradients while incorporating momentum. Although the standard RMSProp does not feature momentum, we found that adding momentum to RMSProp can increase both the performance of DQN and LRRL, aligning with findings in the literature [30, 4]. In the following experiment, we compare the performance of RMSProp with Nesterov’s momentum (RMSProp-M), and Adam when coupled with either LRRL or the best-performing single learning rate when using DQN. Since the results were consistent in previous sections, we extend the experiments to other 3 Atari games.
Figure 4 shows that LRRL coupled with Adam consistently outperforms our configuration using RMSProp-M and the baseline using standard DQN. Moreover, LRRL (RMSProp-M) underperforms compared to DQN without LRRL in two out of three tasks due to its slow convergence despite better jumpstart performance. Future work should investigate whether this slow convergence is due to some optimizer features such as the absence of bias correction in the first and second moment estimates.
Figure 4: A comparison between Adam and RMSProp with momentum, using either DQN or LRRL.
4.4 LRRL with a Distributional RL Algorithm
Lastly, we illustrate the versatility of LRRL by combining it with the distributional RL algorithm Implicit Quantile Networks (IQN) [11], which exhibits the best performance in many Atari games among the algorithms available in Dopamine. Adam is used as the optimizer. Figure 5 shows that LRRL outperforms standard learning with a single learning rate across all the tasks evaluated. More generally, LRRL can be seamlessly integrated with any RL algorithm that employs function approximators based on SGD, improving its performance and mitigating hyperparameter tuning.
Figure 5: Comparison between LRRL and standard IQN learning: IQN combined with LRRL using 3 different learning rates is tested against the 3 variants of standard IQN using each possible learning rate.
@@ 4.5 SGD and Stationary Non-convex Loss Landscapes
Although RL tasks are a challenging showcase to illustrate the effectiveness of our approach in the non-stationary case, there are many confounding factors to analyze their impact on the results, such as the optimizer adaptive routine, the environment stochasticity, and the reward sparseness. To disentangle these factors from our evaluation, in the following experiments we employ the SGD optimizer, which has a single scalar learning rate without adaptive scheme or momentum to find the global minima of $6$ non-convex loss landscapes. Since these functions represent a stationary case, we employ the stochastic bandit Minimax Optimal Strategy in the Stochastic case (MOSS) algorithm (see details in Appendix B.2), using as feedback the loss plus a small constant $\xi$, such that $f_{n}=\frac{1}{|\textrm{loss}(n)|+\xi}$, observed at each gradient step $n$ to update the arm’s probabilities.
Figure 6 illustrates the performance of LRRL against SGD with a single learning rate, starting from the same initial point and for the same amount of steps. The final steps are indicated by a star icon with a different color. The results show that LRRL, with its “all-in-one” strategy, efficiently navigates intricate landscapes. It is effective in combining higher learning rates to accelerate convergence and lower ones for more fine-grained optimization steps when needed, as evidenced in the loss landscapes depicted in Figures 6(a)6(b), and 6(c). In addition, Figures 6(d)6(e), and 6(f) illustrate that LRRL can be also effective in converging towards the best-performing learning rate within the arms, showcased by its steady convergence towards the global minima.
Figure 6: Comparison between LRRL and fixed learning rates in a stationary setting with SGD. LRRL with 3 learning rates is tested against the 3 constant rates for 6 non-convex losses (shown in Figure 10 in Appendix B.4). LRRL is effective in either converging to the best performing rate or combining the set to achieve better results.
@@ 5 Conclusion
In this work, we introduced LRRL, a meta-learning approach for selecting optimizers’ learning rates on the fly. In any optimization task, the best-performing learning rate or rate scheduler is typically unknown, requiring multiple runs to test each candidate in a wasteful computational effort. LRRL mitigates this cost by employing a multi-armed bandit algorithm to explore a set of learning rates within a single run, significantly reducing the need for extensive hyperparameter tuning while achieving better or competitive results with minimal computational overhead. Using RL and non-convex optimization tasks, our empirical results show that combining LRRL with the Adam or SGD optimizers could enhance the performance, outperforming baselines and learning rate schedulers.
Future investigations could extend LRRL ideas to other hyperparameters, such as mini-batch size, which also plays a key role in convergence. Moreover, although LRRL selects rates based on policy performance, alternative feedback mechanisms could be explored, such as using gradient information to select block-wise (e.g., per-layer) learning rates, extending these ideas to supervised learning and applying them to other non-stationary objective functions, including those encountered in Continual Learning [33, 19, 1].
Limitations. The primary limitation of LRRL lies in the need to define a set of arms for the MAB algorithm, as the choice of this set can directly influence the performance of the underlying learning algorithm. However, in computationally intensive settings such as RL tasks, LRRL offers a significant advantage: it enables the simultaneous evaluation of multiple learning rates within a single training run, eliminating the need to test each rate individually. By employing an “all-in-one” strategy, LRRL not only achieves competitive results while reducing the burden of hyperparameter optimization but also can outperform standard learning by combining multiple learning rates.
@@ Appendix A Related Work
This section presents works closely related to ours. We first introduce approaches that use MAB for hyperparameter selection and later other methods that aim to adapt the learning rate over the training process.
@@
Deep RL is known to be overly optimistic in the face of uncertainty [28], and many works have addressed this issue by proposing conservative policy updates [42, 14, 2]. However, when the agent can interact with the environment, this optimism might encourage exploration, potentially leading to the discovery of higher returns. Building on this idea, Moskovitz et al. [27] proposes Tactical Optimism and Pessimism (TOP), an approach that uses an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process. Although our work shares similarities with TOP, such as employing an adversarial bandit algorithm and using performance improvement as feedback, there are key differences. First, TOP is closely tied to actor-critic methods, making it unclear how it could be extended to other RL paradigms. Moreover, it also introduces an exploration strategy which depends on epistemic uncertainty (captured through ensembles) and aleatoric uncertainty (estimated using a distributional RL algorithm). Second, TOP trades-off between optimism and pessimism by adding new variables (arms), which are not inherently tied to the learning process and also require tuning. In contrast, LRRL offers an algorithm-agnostic solution by directly optimizing the learning rate, a core hyperparameter in all SGD-based methods, making it simpler and more broadly applicable. Also close to our work, Liu et al. [24] propose Adam with Bandit Sampling (ADAMBS), which employs the Exp3 algorithm to enhance sample efficiency by incorporating importance sampling within the Adam optimizer. ADAMBS prioritizes informative samples by keeping a probability distribution over all samples, although the feedback is provided by the gradient computed over a mini-batch, unlike previous works in RL that are capable of assigning per-sample importance by using their computed TD-error (see 34).
@@
In order to select the learning rate for stochastic gradient MCMC, Coullon et al. [10] employ an algorithm based on Successive Halving [17, 16], a MAB strategy that promotes promising arms and prunes suboptimal ones over time. In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [22, 36]. A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate on the fly from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space. Close to our work, Parker-Holder et al. [29] propose population-based training which employs a Gaussian Process (see details in 31) with a time-varying kernel to represent the upper confidence bound in a bandit setting. The idea is to have many instances of agents learning a task in parallel with different hyperparameters, replacing those that are underperforming within a single run.
@@
Optimizers such as RMSProp [41] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [18] also incorporates momentum to smooth gradient steps. However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see 13). Increment-Delta-Bar-Delta (IDBD), introduced by Sutton [38], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [47]. Learning rate schedulers with time decay [35, 46] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning. A meta-gradient RL method is proposed in Xu et al. [45], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function. Our work differs from these methods by employing a MAB approach to dynamically select the learning rate over the training process, specifically targeting RL settings.

### v3

@@ 4 Experiments
In this section, we evaluate LRRL across both discrete and continuous control tasks, using multiple RL algorithms and optimizers. Our study includes Atari games (discrete control, DQN and IQN agents) with baselines provided by the Dopamine framework [10], and MuJoCo tasks (continuous control, PPO agents) with baselines implemented in TorchRL [8]. For Atari, we report the mean and one-half standard deviation of returns over 5 independent seeds, while for MuJoCo we use 10 seeds per algorithm. Across all tasks, we compare LRRL against fixed learning rates and schedulers, combined with widely used optimizers such as Adam and RMSProp.
As testing the full Atari benchmark is computationally prohibitive, we select a representative subset of games guided by their baseline learning curves, capturing a range of properties such as reward sparsity, stochasticity, and exploration difficulty. To further disentangle RL-specific confounders, such as environment stochasticity and exploration dynamics, we also evaluate LRRL with SGD on stationary non-convex optimization problems, providing a controlled setting where performance can be attributed solely to learning rate adaptation. Additional details on evaluation metrics and hyperparameters are provided in Appendices D and E.
4.1 LRRL vs. Fixed Learning Rates: Robustness Across Candidate Sets
In our first experiment, we evaluate whether LRRL can remain competitive with tuned baselines when learning rates are drawn from different candidate sets. We consider five possible learning rates,
@@
and construct multiple LRRL variants based on different subsets: the full set $\mathcal{K}(5)$, the three smallest rates $\mathcal{K}{\text{lowest}}(3)$, the three middle rates $\mathcal{K}{\text{middle}}(3)$, the three largest rates $\mathcal{K}{\text{highest}}(3)$, and a sparse mix of low, medium, and high values $\mathcal{K}{\text{sparse}}(3)$. All experiments use the Adam optimizer.
Figure 1: LRRL and standard DQN comparison: 5 variants of LRRL with different learning rate sets are tested against the DQN algorithm reaching best performance among the possible learning rates (provided in Appendix 7). The mean and one-half standard deviations over 5 runs are represented.
It is important to note that the performance of LRRL naturally depends on the composition of the candidate set. If most of the available learning rates perform poorly in isolation, then LRRL’s performance will also be limited, as it can only adapt within that set. This does not indicate a failure of the method, but rather reflects the quality of the available candidates. Conversely, when the set contains both effective and ineffective rates, LRRL consistently adapts toward the better-performing ones, maintaining stability even when poor candidates are present.
In this experiment, we compare LRRL against the best single fixed rate among the candidates. This baseline represents an oracle that assumes prior knowledge of the best-performing learning rate, requiring multiple training runs. LRRL, by contrast, adapts within a single run.
To better understand how this adaptation occurs, we analyze the sequence of learning rates selected during training. Figure 2 visualizes the arms pulled by LRRL $\mathcal{K}(5)$ and the corresponding returns from a single run. In most environments, LRRL initially favors higher learning rates to accelerate early progress, before gradually shifting toward smaller ones as performance stabilizes. This behavior resembles a decay schedule, but crucially, it emerges from policy performance feedback rather than being predefined. The ability to combine aggressive and conservative updates within a single trajectory illustrates how LRRL dynamically adapts to non-stationary learning dynamics.
@@
4.2 LRRL with Learning Rate Schedulers
@@
We next investigate whether LRRL can improve upon fixed learning rate schedulers by adapting among them during training. Specifically, we consider exponential decay schedules of the form $\eta(n)=\eta_{0}\times e^{-dn}$, where $\eta_{0}=6.25\times 10^{-5}$ is a fixed initial value, $d$ is the exponential decay rate, and $n$ is the number of policy updates (i.e., MAB rounds). We define a candidate set $\mathcal{K}_{s}$ of three schedulers with $d\in{1,2,3}\times 10^{-7}$, and compare LRRL against each scheduler used individually, with Adam as the optimizer.
Figure 3 shows that LRRL achieves clear gains in two out of six tasks and remains competitive in the others. Importantly, LRRL consistently improves jumpstart performance in most of tasks. This highlights a key advantage over fixed schedulers: whereas decay schedules impose a monotonic reduction in learning rate, LRRL adapts the choice dynamically based on policy performance. For reference, the dashed black line indicates the performance of Adam with a constant learning rate, which is generally slightly weaker than the best-performing decayed scheduler, in line with previous findings [3].
@@
Figure 3: LRRL vs 3 individual schedulers ($d=1,2,3$). The dashed black line represents the max average return achieved by Adam with a constant learning rate set to $6.25\times 10^{-5}$.
Effect of the Update Window $\kappa$. In the experiments so far, we have fixed the update window $\kappa$ to $1$, meaning LRRL observes a complete episode before updating the learning rate. However, in highly stochastic environments, updating after every episode may provide noisy feedback, making it harder for LRRL to identify the most effective learning rate. Increasing $\kappa$ allows more interactions before each update, trading off reactivity for stability. Two factors should guide this choice:
  • Network update rate: If learning rates are switched more frequently than model updates occur, the feedback to the bandit become inaccurate.
  • Environment feedback: In tasks where outcomes are only revealed after long action sequences, observing more steps before updating helps capture agent performance more reliably.
We evaluate LRRL with different values of $\kappa$ when combined with schedulers. As shown in Figure 4, the effect of $\kappa$ is task-dependent: some games benefit from more reactive updates, while others gain from longer observation windows. Overall, $\kappa=1$ remains a robust choice for Atari games, though higher values can yield slight improvements in certain tasks.
4.3 Continuous Action Spaces
In the following, we evaluate LRRL on MuJoCo tasks using Proximal Policy Optimization (PPO) [36] implemented in TorchRL [8], which by default applies a linear decay schedule to both actor and critic learning rates.
We compare LRRL against two alternative strategies: AdamRel [15] and Cyclical Learning Rates (CLR) (see details in B.5), the latter of which has shown promising results in RL despite higher instability [26]. To isolate the effect of the adaptation mechanism (rather than LR range), we adopt the following implementation choices for this comparison: (i) CLR cycles deterministically over the same candidate learning rate set used by LRRL; (ii) AdamRel uses the standard fixed learning rate (no decay) in this experiment [^1] AdamRel could be combined with decay; CLR could cycle over other ranges. We fix these choices to focus on the adaptation criterion.; and (iii) actor and critic share the same learning rate across methods. We report means and one-half standard deviations over 10 seeds, evaluating policies after each batch for a total of 1M frames.
As shown in Figure 5, LRRL achieves higher returns on two of five MuJoCo tasks. Unlike CLR, which cycles through rates without a performance-based criterion, LRRL adapts its selection from the same pool using policy feedback, leading to stronger or comparable outcomes in most of the tasks. Compared to AdamRel, LRRL does not rely on optimizer-specific mechanics (momentum resets) and is applicable across different optimizers while still adapting to non-stationary learning dynamics.
Figure 5: Comparison of LRRL, AdamRel, and CLR on MuJoCo tasks using PPO. LRRL achieves higher average returns in two tasks. Lines represent the mean and shaded regions show one-half standard deviations over 10 seeds, with training run for 1M frames.
@@ 4.4 SGD and Stationary Non-convex Loss Landscapes
Although RL tasks are a natural showcase for LRRL under non-stationarity, they involve many confounding factors, such as the adaptive routine of the optimizer, stochastic environment dynamics, and reward sparsity. To disentangle these effects, we evaluate LRRL in a simpler, stationary setting: minimizing six standard non-convex loss landscapes using stochastic gradient descent (SGD). Unlike Adam or RMSProp, SGD employs a single scalar learning rate without momentum or adaptivity, making it an ideal testbed to isolate the impact of dynamic learning rate selection.
Experimental setup.
Each minimization works for a fixed budget of steps from the same initialization. LRRL is applied on top of SGD using the stochastic bandit Minimax Optimal Strategy in the Stochastic case (MOSS; see Appendix B.2), with arms corresponding to candidate learning rates. At each gradient step $n$, the bandit receives the feedback
$f_{n}=\frac{1}{\mathcal{J}(w_{1}^{(n)},w_{2}^{(n)})+\xi},\quad\xi>0,$
where $\mathcal{J}$ denotes the loss and $\xi$ prevents division by zero. The bandit updates after every gradient step. Figure 6 illustrates the optimization trajectories, with final iterates marked by stars. Two consistent behaviors emerge:
  • Adapting across multiple rates. On landscapes such as Beale, Bohachevsky, and Griewank, LRRL flexibly alternates between larger and smaller learning rates depending on the stage of optimization. By “playing” with multiple rates, LRRL can accelerate progress when higher values are beneficial and switch to smaller ones when finer convergence is required. This adaptability allows it to escape poor local minima and achieve stronger convergence than any single fixed rate.
  • Converging to the best candidate. On functions where a single rate is clearly optimal (e.g., Rosenbrock, Zakharov), LRRL adapts toward that rate, outperforming the second best performing candidate.
These experiments demonstrate that LRRL is not only robust to candidate quality but can also exploit multiple learning rates within a single run. Whereas a fixed learning rate must trade off between speed of convergence and stability, LRRL adapts dynamically, combining both advantages. Importantly, these results in stationary optimization confirm that LRRL’s effectiveness is not limited to the non-stationary structure of RL, but reflects a more general capacity to adapt learning dynamics.
Figure 6: LRRL vs fixed learning rates in a stationary setting with SGD. LRRL with 3 rates is tested against the 3 constant rates for 6 non-convex losses (shown in Figure 11 in Appendix). LRRL is effective in either converging to the best performing rate or combining the set to achieve better results.
@@ 5 Conclusion
We introduced LRRL, a meta-learning approach that adapts the optimizer’s learning rate on the fly using a multi-armed bandit. In contrast to fixed schedules or oracle tuning that require multiple runs, LRRL evaluates and selects among candidate rates within a single run, reducing tuning effort while remaining competitive or superior to the best fixed choice. Across discrete and continuous RL tasks, LRRL improved or matched strong baselines such as decay schedules, AdamRel, and CLR, and generalized to stationary non-convex optimization with SGD. These results highlight LRRL’s ability to combine aggressive and conservative updates within a single trajectory, offering both robustness and adaptability.
Limitations. LRRL operates over a predefined set of candidate learning rates, which serves as the search space for adaptation. However, this requirement is inherent to any hyperparameter search, and our experiments show that LRRL can exploit multiple candidates dynamically, avoiding catastrophic failures and often outperforming fixed-rate baselines. Theoretical guarantees would require strong assumptions (e.g., convexity, smoothness, bounded losses) that deviate from deep RL; instead, we focused on broad empirical validation.
@@ Appendix A Related Work
This section presents works closely related to ours. We first introduce approaches that use MAB for hyperparameter selection, and later other methods that aim to adapt the learning rate over the training process.
@@
Deep RL is known to be overly optimistic in the face of uncertainty [31], and many works have addressed this issue by proposing conservative policy updates [46, 16, 1]. However, when the agent can interact with the environment, this optimism might encourage exploration, potentially leading to the discovery of higher returns. Building on this idea, Moskovitz et al. [30] proposes Tactical Optimism and Pessimism (TOP), an approach that uses an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process. Although our work shares similarities with TOP, such as employing an adversarial bandit algorithm and using performance improvement as feedback, there are key differences. First, TOP is closely tied to actor-critic methods, making it unclear how it could be extended to other RL paradigms. Moreover, it also introduces an exploration strategy which depends on epistemic uncertainty (captured through ensembles) and aleatoric uncertainty (estimated using a distributional RL algorithm). Second, TOP trades-off optimism and pessimism by adding new variables (arms), which are not inherently tied to the learning process and also require tuning. In contrast, LRRL offers an algorithm-agnostic solution by directly optimizing the learning rate, a core hyperparameter in all SGD-based methods, making it simpler and more broadly applicable. Also close to our work, Liu et al. [24] propose Adam with Bandit Sampling (ADAMBS), which employs the Exp3 algorithm to enhance sample efficiency by incorporating importance sampling within the Adam optimizer. ADAMBS prioritizes informative samples by keeping a probability distribution over all samples, although the feedback is provided by the gradient computed over a mini-batch, unlike previous works in RL that are capable of assigning per-sample importance by using their computed TD-error (see 35).
@@
In order to select the learning rate for stochastic gradient MCMC, Coullon et al. [11] employs an algorithm based on Successive Halving [19, 18], a MAB strategy that promotes promising arms and prunes suboptimal ones over time. In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [23, 38]. A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate on the fly from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space. Parker-Holder et al. [32] propose population-based training which employs a Gaussian Process (see details in 34) with a time-varying kernel to represent the upper confidence bound in a bandit setting. The idea is to have multiple instances of agents learning a task in parallel, each with different hyperparameters, and replace those that are underperforming within a single run. Close to our work, MABSearch [43] frames the learning rate selection as a MAB. In practice, however, its algorithm reduces to an heuristic $\epsilon$-greedy scheme that maintains an exponential moving average of the loss to guide exploitation. As such, its performance depends critically on the choice of $\epsilon$ and its decay, lacking the principled, adaptive exploration–exploitation balance over the long run that defines proper bandit methods. In contrast, LRRL builds on adversarial algorithms such as Exp3, which maintain probabilistic weights over arms and update them under non-stationary feedback, providing the first rigorous integration of bandits as a meta-optimizer for learning rate adaptation.
@@
Optimizers such as RMSProp [45] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [20] also incorporates momentum to smooth gradient steps. However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see 14). Increment-Delta-Bar-Delta (IDBD), introduced by Sutton [41], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [51]. Learning rate schedulers with time decay [37, 50] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning. A meta-gradient RL method is proposed in Xu et al. [49], formulated as a two-level optimization process: one level optimizes the agent’s policy objective, while the other learns meta-parameters of the objective function. Although both LRRL and meta-gradient approaches aim to enable agents to “learn how to learn” within a single agent’s lifetime, meta-gradient methods must be tailored to the specific loss structure and update rules of each RL algorithm. In contrast, LRRL is algorithm-agnostic: it operates at the optimizer level, adapting the learning rate through a bandit mechanism without requiring gradients of the learning objective or modifications according to the underlying algorithm.